# Scalable Stewardship of an LLM-Assisted Clinical Benchmark with Physician Oversight

### v2

@@ Abstract
Background: Medical scores, such as Glasgow Coma Scale and Charlson Comorbidity Index, translate textual and tabular patient data into standardized risk metrics that augment triage, treatment, and prognostication workflows for direct patient care. Traditionally, these scores are manually evaluated by clinicians, and their repetitiveness imposes a major clerical burden. MedCalc-Bench recently introduces a large-scale clinical benchmark to evaluate how effective a Large Language Model (LLM) could automate 55 common calculation tasks. However, to reduce the labor intensity of manual annotations, MedCalc-Bench itself used LLMs (GPT-3.5 and GPT-4) to derive the reference score labels, which could embed historical model biases. This study aims to assess how severely these reference labels are clinically inaccurate and to identify scalable approaches for improving label quality.
Methods: We developed a phased stewardship process to audit and improve MedCalc-Bench labels at scale. The process begins with two separate, knowledge-grounded LLM agentic workflows: an audit of the original test cases including their labels’ derivation metadata, and an independent relabeling of the original train and test cases from first principles. This automated triage system spotlights cases with high-disagreement between original and recomputed reference labels, to help concentrate a finite budget of physician experts’ feedback on the most contentious instances. We also quantified the impact of uncaught label noise through a controlled reinforcement-learning (RL) study, comparing models trained on original versus our recomputed reference labels to assess performance impacts.
Results: On the MedCalc-Bench test set (n=1,047), the audit flagged the labels of 279 cases (26.6%) as clinically suspect, spanning 40 of 55 calculation tasks. A clinician spot-check agreed with its error identification in 7/7 sampled cases. In a separate independent recomputation, at least 286/1,047 original test labels (27.3%) either exceeded the benchmark’s tolerance relative to recomputed values or were deemed incomputable. In an alignment study (n=50), recomputed labels agreed with licensed physicians’ gold standard labels more often than original labels (37/50 vs 10/50) and substantially reduced sMAPE error from 72.7% (95% CI, 47.9-99.1) to 20.1% (95% CI, 8.0-35.9). Lastly, in the controlled RL experiment, updating the training labels shifted measured test accuracy (71.4% vs 62.6%), illustrating how label quality can materially affect downstream conclusions about model performance.
Conclusion: Physician-in-the-loop stewardship is essential to prevent clinical benchmarks from becoming “biased teachers.” Our results demonstrate that partially LLM-generated labels can embed systemic errors that distort not only evaluation but also downstream model alignment. By adopting a hybrid oversight system, we can prioritize scarce expert feedback to maintain benchmarks as living, clinically-grounded documents. Ensuring this alignment is a prerequisite for the safe deployment of LLMs in high-stakes medical decision support.
Short Description
We examine the reliability of a widely used clinical AI benchmark whose reference labels were partially generated by LLMs, and find that a substantial fraction are clinically misaligned. We introduce a phased stewardship procedure to amplify the positive impact of physician experts’ feedback and then demonstrate, via a controlled RL experiment, how uncaught label bias can materially affect downstream LLM evaluation and alignment.
@@ 1 Introduction
A clinician’s day often alternates between patient care, time spent on complex judgment, and time spent on repetitive “information work”: retrieving scattered facts from the electronic health record (EHR), reconciling fragmented documentation, and translating that information into standardized actions. This clerical burden contributes to burnout and has been linked to patient-safety and health-system consequences.[14, 37, 27, 39] Tools based on Large Language Models (LLM) are increasingly proposed to reduce this burden, including ambient documentation support and targeted safety guardrails in healthcare delivery.[40, 1, 30, 18, 8]
Routine computation of medical risk and severity scores is another common information processing scenario that appears suitable for automation. Risk scores such as LACE, CURB-65, and CHA2DS2-VASc condense patient history, vital signs, and laboratory results into numeric values that inform triage, admission planning, and preventive therapy;[42, 20, 22] a worked out example in Supplementary Appendix §D.1 illustrates a computation routine. In practice, clinicians extract criteria from narrative notes or structured EHR data, and then manually calculate the final scores or enter information into calculators such as MDCalc.[26] Manual extraction is time-consuming and error-prone, motivating the adoption of large language model (LLM) systems that can read notes and compute scores end-to-end.
As these automated systems are developed, evaluating their efficacy in a scalable but also clinically aligned manner becomes a safety challenge: how do we responsibly verify that they compute scores as accurately as would a conscientious human physician? Public task benchmarks are an important first step for transparency and reproducibility, but large-scale benchmark construction is increasingly assisted by LLMs themselves through feature extraction and normalization. As a prime example, MedCalc-Bench [16] is a widely used large-scale public benchmark for end-to-end medical risk score computation. Besides popularity in academic research,[6, 23, 21, 47, 43, 28] it was one of two medical benchmarks evaluated on by Anthropic in its recent launch of the “Claude for Healthcare” initiative.[3, 4]
MedCalc-Bench was constructed from de-identified PubMed Central case reports paired with 55 commonly-used medical calculators on MDCalc. Its original labeling pipeline used GPT-4 for feature extraction and scripted aggregation for final score computation. This automated labeling pipeline yielded a large task dataset of 11,000 labeled instances. The scale and comprehensive calculator coverage of MedCalc-Bench make it an important contribution to clinical natural language processing (NLP) research, as AI model training is inherently data-hungry. Manually producing benchmark labels at this scale is difficult, as sample-by-sample clinician adjudication is slower and much more expensive than LLM API calls: sizes of fully physician-labeled medical score datasets, such as [11] and [32], are in the scale of low hundreds. However, the expediency introduces a structural vulnerability: labels can inherit systemic failure modes of the language models, tools, and procedural assumptions used to produce reference labels. It is a concerning situation because in effect, these synthesized benchmark labels are not only the “ground truth” used to measure progress and, increasingly, also a target distribution for AI developers to align their models towards, for example via reinforcement learning (RL).[29, 2] When the same labels are reused as RL reward signals, label errors are no longer only measurement noise: they can actively shape the behavior of future models, turning the benchmark from a miscalibrated yardstick into a biased teacher.
In this paper, we audit and improve the clinical soundness of reference labels originally released by MedCalc-Bench in its NeurIPS 2024 publication.[16] Schematically, we propose a scalable stewardship protocol in which agentic LLM pipelines—LLM-based verifiers instructed to follow a structured checklist and use tools such as search and computation—audit existing labels and independently recompute scores; automated triage then concentrates scarce physician time on the highest-impact disagreements. Physician review resolves contention and identifies truly unanswerable or mismatched note-question pairs, for which abstention (e.g., “unknown”) is safer than forced computation. Importantly, our goal is not to fault the creators of MedCalc-Bench; rather, we use their original dataset release as a motivating case study to develop a transparent, repeatable stewardship process for benchmarks that increasingly influence both evaluation and post-training. Our contributions are as follows:
  • Propose systematic stewardship for clinical workflow benchmarks as safety infrastructure and describe the risks of treating partially LLM-generated labels as static oracles.
  • Introduce a physician-in-the-loop protocol that combines agentic audit, independent recomputation, and automated triage, with clinician adjudication for disputed cases.
  • Apply the protocol to MedCalc-Bench, characterize major sources of label failure, and incorporate an abstention label for unanswerable note-question pairs.
  • Demonstrate in a controlled RL setting that correcting for latent training label bias can materially change LLM behavior and performance interpretation.
The remainder of the manuscript describes the benchmark’s original label-generation process and label failure modes, presents our stewardship protocol and physician review process, and then examines how revised labels change evaluation and post-training conclusions.
@@ 2.1 Benchmark setting and task definition
MedCalc-Bench is a public benchmark designed to evaluate whether language models can compute common medical scores from free-text narratives, without requiring external calculator use.[16] Each benchmark instance presents a model with a clinical context, typically several paragraphs of a de-identified patient narrative, and a question requesting a specific medical score (e.g., “What is the patient’s HAS-BLED score?”). The model must return either a numeric value, a datetime, or, in our extended formulation, an explicit abstention ($\mathsf{N/A}$”) when the context lacks necessary information or is clinically mismatched to the question. The benchmark contains 11,000 context-question pairs (10,053 training; 1,047 test) covering a set of 55 different score questions. This task mirrors what clinicians do: read documentation, identify the criteria a particular score requires, and apply the scoring rule. The benchmark therefore tests whether a model can perform both subtasks, criterion extraction from unstructured text and rule-based aggregation, correctly and in sequence.
@@
Our analysis focuses on the peer-reviewed release of MedCalc-Bench (v1.0), which serves as the standard baseline for current academic and industry evaluations (e.g., [6], [3]; see also the last paragraph of Appendix §D.2). We note that while the authors began releasing updates (v1.1, v1.2) during the final preparation of this manuscript, our auditing and reinforcement learning experiments were completed prior to these releases. We therefore utilize v1.0 to ensure: (1) comparability with the existing body of concurrent work, and (2) that our stewardship protocol is tested against the distribution of errors present in the widely circulated version of the benchmark. A detailed analysis of the performance discrepancies between dataset versions is provided in Supplementary Appendices B and G.
@@
In MedCalc-Bench’s publication[16], reference labels were produced by a three-stage procedure: GPT-3.5 matched patient contexts to appropriate calculator questions, GPT-4 extracted clinical features from each narrative (e.g., patient demographics, symptoms, laboratory values), and custom Python scripts aggregated those features according to each calculator’s formula or decision logic. At test time, a model prediction is scored correct if it falls within $\pm$5% of the reference label for numeric outputs or matches exactly for datetimes. We extend this scheme to treat abstention as correct when our recomputed reference label is itself “unknown.”
@@
We define physician-aligned ground truth, $y^{\star}$, as the answer a conscientious physician would provide for a given note-question instance $(C,q)$. The corresponding benchmark reference label, which we denote by $\hat{y}$, is an estimate of this (latent) ground truth; $\hat{y}^{\mathrm{original}}$ is the reference label originally given by MedCalc-Bench, while $\hat{y}^{\mathrm{new}}$ is the label our agentic workflow independently recomputes in §2.4. Estimation error can enter at any of the three stages, whether through inappropriate context-calculator pairing, hallucinated or misread feature values, or aggregation scripts that diverge from published scoring rules. Task-level ambiguity, such as missing timepoints or guideline version differences, introduces additional uncertainty. Recognizing these labels as fallible estimates, rather than one-off oracles, motivates the stewardship process described below.
@@
Figure 1: Our benchmark stewardship utilizes two distinct LLM agent workflows (Phases 1 & 2) to assure MedCalc-Bench’s label quality. Besides their prompt and output type (Yes/No verdict versus a recomputed label), the two workflows differ in what the LLM agent is given as context: the Phase 1 agent is shown the original reference label and its derivation metadata, whereas the Phase 2 agent is only provided the patient narrative and score question; the latter omission is intended to reduce any anchoring bias towards $\hat{y}^{\mathrm{original}}$. In Phase 3, three physician authors (DT, AG, NK) independently compute 50 instances where $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ show high disagreement. sMAPE stands for symmetric Mean Absolute Percentage Error (Eq. 2); it measures the relative disagreement in $\{(y^{\star}_{i},\hat{y}^{\mathrm{new}}_{i})\}_{i=1}^{50}$ and $\{(y^{\star}_{i},\hat{y}^{\mathrm{original}}_{i})\}_{i=1}^{50}$.
@@ 2.2 Stewardship process overview
The central constraint in validating benchmark reference labels is not machine computation but physician experts’ attention. GPU hours are elastic; physician hours are not. Our process is therefore designed around a principle of phased resource commitment: each phase produces evidence that justifies, or gates, incremental investment, concentrating scarce expert time on instances most likely to benefit from it. This design balances scalability through LLM-based meta-evaluation with the authority of physician judgment, which remains the ultimate source of clinical ground truth but is routed to high-information decisions.
The process proceeds in three phases, in which the first two serve as an automated instances triage. Phase 1 (automated audit) applies a tool-augmented LLM verifier to existing labels, flagging instances where the original derivation of $\hat{y}^{\mathrm{original}}$ appears clinically unsound. This phase is inexpensive to run and does not commit to replacement labels; its purpose is to provide an early signal of the overall error rate, justifying further investigation. Phase 2 (independent recomputation) deploys a separate LLM pipeline that sees only the patient context and question, not the original label or its provenance, and recomputes a label $\hat{y}^{\mathrm{new}}$ from first principles. Disagreement between the recomputed label and the original, beyond a clinically meaningful threshold, identifies cases where at least one estimate is wrong. Finally, Phase 3 (physician adjudication) directs clinician review to the subset of instances where Phase 2 recomputation diverges most substantially from original labels, maximizing information gained per expert hour.
@@ 2.3 Phase 1: Tool-augmented automated audit of original labels
We implemented the auditor as a reasoning language model (Gemini 2.5 Pro, June 2025 version) configured as a verifier with web-search access. During each audit, the agent could retrieve calculator definitions, rule tables, and guideline statements as needed, improving factual grounding beyond parametric knowledge.
@@
We use a sampling temperature of 1 for generating outputs, following Google’s official recommendation.[12] Because LLM outputs are stochastic, we ran the auditing agent five times independently for each $(C,q,\hat{y}^{\mathrm{original}})$ sample using identical instructions. We then aggregated verdicts with a conservative supermajority rule: an instance was labeled Likely Error only if at least 4 of 5 runs flagged $\hat{y}^{\mathrm{original}}$ as incorrect in a clinically significant way; otherwise it was labeled Passed. This design is intended to prioritize precision over recall.

### v3

@@ Abstract
Reference labels for machine-learning benchmarks are increasingly synthesized with LLM assistance, but their reliability remains underexamined. We audit MedCalc-Bench, a clinical benchmark for medical score computation whose labels were partly derived with LLM assistance, and develop a scalable physician-in-the-loop stewardship pipeline to reassess them. At least 27% of test labels are likely erroneous or incomputable. On a 50-instance subset validated by physicians, our recomputed labels agree with physician ground truth 74% of the time (95% CI, 60–84%) versus 20% for the originals (95% CI, 11–33%). Using original labels to evaluate frontier LLMs underestimates accuracy by 16–23 percentage points. In a controlled reinforcement-learning experiment, a model trained on recomputed labels outperforms one trained on originals by 13.5 percentage points (95% CI, 10.6–16.6%) on physician-labeled instances, and this advantage extends to related medical tasks. LLM-assisted benchmarks can propagate systematic errors into both evaluation and post-training unless actively stewarded.
@@ 1 Introduction
Benchmarks have become the primary infrastructure for measuring AI capabilities. The standard recipe pairs a set of tasks with reference answers and a verification procedure, producing a score that summarizes model performance on a given capability (Bedi et al., 2026). As large language models (LLMs) are deployed in higher-stakes settings, the demand for domain-specific benchmarks has grown accordingly: in medicine alone, recent work has introduced benchmarks for clinical question answering, diagnostic reasoning, and risk-score computation (Jin et al., 2021; Khandekar et al., 2024; Goodell et al., 2025). These benchmarks shape not only how models are evaluated but increasingly how they are trained, as benchmark labels are reused as reward signals for reinforcement learning (RL) and other post-training methods (Ouyang et al., 2022; Ahmadian et al., 2024).
Building benchmarks at sufficient scale is itself a major challenge. Statistically meaningful evaluation requires large sample sizes, yet producing reference labels in specialized domains demands expensive expert labor: in medicine, each label may require a physician to review the patient’s chart and exert clinical judgment to arrive at an answer. Fully physician-labeled datasets therefore tend to be small (Goodell et al., 2025; Roeschl et al., 2025). This tension between scale and quality has made it natural, even necessary, to enlist LLMs in the benchmark-creation process itself. The result is a circularity: AI systems help build the very evaluation infrastructure used to judge AI systems. Label errors in benchmarks are not hypothetical: prior work has shown that commonly used ML datasets harbor pervasive mislabeling that can destabilize model rankings (Northcutt et al., 2021), and recent studies have used LLM ensembles to detect and correct such errors in NLP benchmarks (Nahum et al., 2024). However, in high-stakes domains such as medicine, where expert labeling costs are high and the consequences of benchmark error are particularly severe, systematic evidence remains scarce.
We provide such evidence in one of the high-stakes domains. Medical risk scores such as LACE, CURB-65, and CHA2DS2-VASc condense patient data into numeric values that inform triage, admission, and preventive therapy (Van Walraven et al., 2010; Lim et al., 2003; Lip et al., 2010). Automating their computation from clinical narratives is a natural target for LLMs (Tierney et al., 2024; Afshar et al., 2025; Pais et al., 2024; Korom et al., 2025; Chung et al., 2025), and MedCalc-Bench (Khandekar et al., 2024) is the most widely used public benchmark for this task. Its labeling pipeline used GPT-4, the leading model at the time of the benchmark’s construction, for feature extraction and scripted aggregation to produce 11,100 labeled instances across 55 calculators. The benchmark has been adopted broadly in academic research (Bedi et al., 2026; Liu et al., 2025a; Lin et al., 2025; Zhang et al., 2025; Wang et al., 2025; Oh et al., 2025) and was one of two medical benchmarks evaluated on by Anthropic in its recent “Claude for Healthcare” launch (Anthropic, 2026a, b). Despite this reach, the clinical accuracy of its LLM-generated labels has not been systematically assessed.
In this paper, we develop a scalable stewardship protocol and apply it to MedCalc-Bench. At its core, multiple tool-augmented LLM agents independently recompute each score from the patient narrative alone, without access to the original label, and abstain when the clinical context is insufficient or mismatched to the question. A consensus mechanism across agents produces high-confidence recomputed labels. Automated triage then identifies instances where the recomputed and original labels diverge most, concentrating scarce physician time on the highest-impact disagreements; physicians independently recompute scores for these contentious cases, providing ground-truth adjudication.
We find that a substantial fraction of MedCalc-Bench test labels are likely erroneous or incomputable. We examine the downstream consequences of these errors on two fronts: first, by re-evaluating today’s frontier LLMs against our corrected labels and comparing with officially reported results, including those from Anthropic’s “Claude for Healthcare” launch (Anthropic, 2026a); and second, through a controlled RL experiment in which the only variable is whether training rewards are derived from the original or corrected labels. Both analyses demonstrate that label quality materially affects conclusions about model capabilities. MedCalc-Bench was a significant undertaking that laid the foundation for studying LLM performance on medical score computation; we view our work as building on that foundation rather than critiquing it, offering a transparent, repeatable stewardship process and arguing that benchmarks in safety-critical domains are best maintained as living documents rather than treated as static artifacts.
@@ 2.1 Benchmark setting and task definition
MedCalc-Bench is a public benchmark designed to evaluate whether language models can compute common medical scores from free-text narratives, without requiring external calculator use (Khandekar et al., 2024) (a worked example using the LACE score is provided in the Results section and Table 1 in Supplementary Appendix §D). Each benchmark instance presents a model with a clinical context, typically several paragraphs of a de-identified patient narrative, and a question requesting a specific medical score (e.g., “What is the patient’s HAS-BLED score?”). The model must return either a numeric value, a datetime, or, in our extended formulation, an explicit abstention ($\mathsf{N/A}$”) when the context lacks necessary information or is clinically mismatched to the question. The benchmark contains 11,000 context-question pairs (10,053 training; 1,047 test) covering a set of 55 different score questions. This task mirrors what clinicians do: read documentation, identify the criteria a particular score requires, and apply the scoring rule. The benchmark therefore tests whether a model can perform both subtasks, criterion extraction from unstructured text and rule-based aggregation, correctly and in sequence.
@@
Our analysis focuses on the peer-reviewed release of MedCalc-Bench (v1.0), which serves as the standard baseline for current academic and industry evaluations (e.g., (Bedi et al., 2026), (Anthropic, 2026a); see also the last paragraph of Appendix §D.2). We note that while the authors began releasing updates (v1.1, v1.2) during the final preparation of this manuscript, our auditing and reinforcement learning experiments were completed prior to these releases. We therefore utilize v1.0 to ensure: (1) comparability with the existing body of concurrent work, and (2) that our stewardship protocol is tested against the distribution of errors present in the widely circulated version of the benchmark. A detailed analysis of the performance discrepancies between dataset versions is provided in Supplementary Appendices B and G.
@@
In MedCalc-Bench’s publication (Khandekar et al., 2024), reference labels were produced by a three-stage procedure: GPT-3.5 matched patient contexts to appropriate calculator questions, GPT-4 extracted clinical features from each narrative (e.g., patient demographics, symptoms, laboratory values), and custom Python scripts aggregated those features according to each calculator’s formula or decision logic. While GPT-4 was among the most capable models available when MedCalc-Bench was created and LLMs have improved substantially since then, even today’s most capable models in agentic settings still hallucinate and make errors on clinical tasks, underscoring the need for systematic verification of LLM-assisted labels regardless of the model used to produce them. At test time, a model prediction is scored correct if it falls within $\pm$5% of the reference label for numeric outputs or matches exactly for datetimes. We extend this scheme to treat abstention as correct when our recomputed reference label is itself “unknown.”
@@
We define physician-aligned ground truth, $y^{\star}$, as the answer a conscientious physician would provide for a given note-question instance $(C,q)$. The corresponding benchmark reference label, which we denote by $\hat{y}$, is an estimate of this (latent) ground truth; $\hat{y}^{\mathrm{original}}$ is the reference label originally given by MedCalc-Bench, while $\hat{y}^{\mathrm{new}}$ is the label our agentic workflow independently recomputes in §2.4. This two-stage structure, feature extraction followed by rule-based aggregation, is formalized in Supplementary Appendix §D.1. Estimation error can enter at either stage, whether through inappropriate context-calculator pairing, hallucinated or misread feature values, or aggregation scripts that diverge from published scoring rules. Task-level ambiguity, such as missing timepoints or guideline version differences, introduces additional uncertainty. Recognizing these labels as fallible estimates, rather than one-off oracles, motivates the stewardship process described below.
@@ 2.2 Stewardship process overview
The central constraint in validating benchmark reference labels is not machine computation but physician experts’ attention. GPU hours are elastic; physician hours are not. Our process is therefore designed around a principle of phased resource commitment: each phase produces evidence that justifies, or gates, incremental investment, concentrating scarce expert time on instances most likely to benefit from it.
The process proceeds in three phases (Figure 1), in which the first two serve as an automated instances triage. Phase 1 (automated audit) applies a tool-augmented LLM verifier to existing labels, flagging instances where the original derivation of $\hat{y}^{\mathrm{original}}$ appears clinically unsound. This phase is inexpensive to run and does not commit to replacement labels; its purpose is to provide an early signal of the overall error rate, justifying further investigation. Phase 2 (independent recomputation) deploys a separate LLM pipeline that sees only the patient context and question, not the original label or its provenance, and recomputes a label $\hat{y}^{\mathrm{new}}$ from first principles. Disagreement between the recomputed label and the original, beyond a clinically meaningful threshold, identifies cases where at least one estimate is wrong. Finally, Phase 3 (physician adjudication) directs clinician review to the subset of instances where Phase 2 recomputation diverges most substantially from original labels, maximizing information gained per expert hour.
Figure 1: Three-phase stewardship workflow. Phase 1 (audit): an LLM verifier reviews each MedCalc-Bench label against its derivation metadata and flags clinically suspect cases. Phase 2 (recomputation): a separate agent recomputes the label from the patient narrative and question alone, producing $\hat{y}^{\mathrm{new}}$; the original label is hidden to prevent anchoring. Phases 1 and 2 each run five LLM agents independently in parallel and aggregate their outputs by supermajority voting to reduce stochastic noise. Phase 3 (adjudication): three physician authors (DT, AG, NK) blind-recompute the 50 instances where $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ disagree most, producing physician ground truth $y^{\star}$. sMAPE measures the relative gap between each label set and $y^{\star}$ (Eq. (2) in §2.5).
@@ 2.3 Phase 1: Tool-augmented automated audit of original labels
We implemented the auditor as a reasoning language model (Gemini 2.5 Pro, June 2025 version) configured as a verifier with web-search access. We selected Gemini 2.5 Pro because at the time Phases 1 and 2 of this study were conducted (June–October 2025), it was among the highest-ranked models on a wide range of public reasoning and knowledge benchmarks, including those covering scientific and clinical domains. The same model is reused as the Phase 2 recomputation agent (§2.4) to keep the labeling pipeline internally consistent. We separately evaluate today’s frontier LLMs (released after our auditing experiments were completed) against the resulting curated labels in §3.3. During each audit, the agent could retrieve calculator definitions, rule tables, and guideline statements as needed, improving factual grounding beyond parametric knowledge.
@@
We use a sampling temperature of 1 for generating outputs, following Google’s official recommendation (Google Cloud, 2025). Because LLM outputs are stochastic, we ran the auditing agent five times independently for each $(C,q,\hat{y}^{\mathrm{original}})$ sample using identical instructions. We then aggregated verdicts with a conservative supermajority rule: an instance was labeled Likely Error only if at least 4 of 5 runs flagged $\hat{y}^{\mathrm{original}}$ as incorrect in a clinically significant way; otherwise it was labeled Passed.